length of a rectangle is 3cm greater then its width. the perimeter is 24 cm. what is the dimensions of the rectangle
Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

length of a rectangle is 3cm greater then its width. the perimeter is 24 cm. what is the dimensions of the rectangle
-----
Let width = W
length = W+3
------------------
Perimeter = 2(width + length)
24 = 2(W + W+3))
12 = 2W + 3
2W = 9
W = 4.5 cm
length = W+3 = 7.5 cm
=============================
Cheers,
Stan H.

Answer by mananth(16949)

width = x
length = x+3
perimeter p = 2*(L+W)
p=2*(x+x+3)
p=2*(2x+3)
p=4x+6
4x+6 =24
4x =24-6
4x=18
divide by 4
4x/4 = 18/4
x=4.5 cm. width
..
Length = x+3
=4.5+3=7.5 cm.